### v1

# Spin-charge separation and resonant valence bond spin liquid in a frustrated doped Mott insulator

Abstract
Anderson’s groundbreaking ideas of resonant valence bond (RVB) liquid and spin-charge separation initiated a transformative shift in modern physics. Extensive implications influenced a broad spectrum of fields, from high-temperature superconductors to quantum computing, and gave birth to key concepts in physics, such as quantum spin liquids, emergent gauge symmetries, topological order, and fractionalisation. Despite extensive efforts to demonstrate the existence of an RVB phase in the Hubbard model, a definitive realisation has proven elusive. Here, we present a concise, realistic, and elegant solution to this longstanding problem by demonstrating analytically that an RVB spin liquid, exhibiting spin-charge separation, emerges as the ground state of doped Mott insulators on corner-sharing tetrahedral lattices with frustrated hopping near half-filling – a manifestation of the counter-Nagaoka effect. We confirm numerically that this result holds for finite-size systems, finite dopant density, and small exchange interactions. While much attention has been devoted to the emergence of new states from geometrically frustrated interactions, our work demonstrates that kinetic energy frustration in doped Mott insulators may be pivotal to stabilise robust, topologically ordered states in real materials.
Inspired by Pauling’s early theory of resonant covalent-bond-sharing in aromatic molecules, such as benzene, and by Bethe’s antiferromagnetic linear chain, Fazekas and Anderson [Anderson1973, Fazekas1974] proposed in the early 1970’s a possible competitor to Néel order in antiferromagnetically coupled spins: a state where spins are valence bond paired into singlets that resonate quantum mechanically amongst different pairing configurations, thus compensating for the partial loss of antiferromagnetic energy – forming a resonant valence bond (RVB) state. The exciting idea gained further momentum with the discovery of high-temperature superconductivity about a decade later, and with Baskaran and Anderson’s prompt proposal [Anderson1987, Anderson1987b, Anderson1997] that it may indeed be underpinned by spin-charge separation [Baskaran1987, Zhu1988, Anderson1987b, Kivelson1987] in a system of strongly correlated electrons in an RVB liquid phase (for a review, see Refs. Anderson2004, Lee2006). The appeal was so great that for decades to follow numerous physicists have been looking for evidence of this suggestion “along a bewildering variety of routes” [Anderson2004].
Aside from its relevance to high-temperature superconductivity, the quest for an RVB liquid state led to several fundamental advances, many of which play a central role in our understanding of modern physics. Of particular import, the notion of spin liquids and its relation to deconfinement in gauge theory have been greatly clarified [Lee2006, Wen2007], and subsequently extended to further areas of research [Moessner2001, Balents2002, Kitaev2003, Wen2003], to encompass new concepts such as topological order and fractionalisation [Knolle2019].
Figure 1: Pictorial illustration of one of the dimer-singlet (dark thin ellipsoids) states with a holon at a given location (light sphere), that participate in the superposition forming the RVB liquid ground state of the $t$-$J$ model, Eq. (1), on a lattice of corner-sharing tetrahedra. The $\pi$-fluxes depict the phase induced by the singlet coverings (relative to a ferromagnetic background) when the holon moves around the corresponding triangular plaquette. The inset shows an Y2Ir2O7 pyrochlore crystal grown by D. Prabhakaran in Oxford.
Intense work continued throughout the years, yet Anderson’s original proposal has hitherto evaded discovery. Following the seminal work by Moessner and Sondhi [Moessner2001], which established the possibility of a short-ranged liquid phase in quantum dimer models, several efforts attempted to translate this result into a system of SU(2) spins, building on Klein models [Klein1982], and effective and decorated dimer models [Batista2004, Raman2005, Normand2014, Normand2016]. It is generally believed that realistic Hamiltonians harbouring RVB spin liquid phases and spin-charge separation must exist, but to date they have remained elusive.
In this work, we find a concise and elegant solution – unintentionally fulfilling the inspirational words of Ref. Raman2005: “a more important task perhaps, now that the question of principle is settled, is to refocus on studying much simpler Hamiltonians”. We demonstrate that an RVB liquid phase, exhibiting spin-charge separation, emerges naturally upon doping a large-$U$ Hubbard model on a lattice of corner-sharing tetrahedra – be it the planar checkerboard or the three-dimensional pyrochlore lattice – near half-filling. This result hinges on kinematic spin correlations set by frustrated hopping and is an instance of the so-called ‘counter-Nagaoka effect’ [Haerter2005]. Without loss of generality, in the following we shall consider the case of hole doping; the same can be said for electron doping when the hopping amplitude has the opposite sign. We prove our claims by finding an exact lower bound to the Hamiltonian energy, and by explicitly constructing an RVB liquid state that precisely matches it in the thermodynamic single-hole limit. We confirm this behaviour by investigating numerically systems of finite size. We propose a way to construct single-hole and two-holes RVB liquid ground state wave functions for any size that appear to be exact within numerical accuracy – a truly remarkable feat for a strongly correlated electron problem. Finally, we show numerically that our results are stable in systems of finite hole density, as well as in presence of sufficiently small exchange interactions between the spins [footnote:Normand].
The crucial role of kinetic energy frustration in stabilising a topologically ordered RVB liquid state is a remarkable outcome of our findings, which provide a novel way to connect dimer coverings with solutions to the Hubbard model. Our results have the potential to shift the traditional paradigm, which has predominantly considered exchange frustration as the mechanism for stabilising quantum spin liquids. In other words, doping insulators that lie deep within the Mott regime with a single type of carrier (electrons or holes, depending on the sign of the hopping amplitude) emerges as a new guiding principle for the discovery of spin liquid states in correlated matter.
Our Hamiltonian is eminently simple and realistic, and of direct relevance to several experimental settings as well as synthetic quantum platforms [Feynman1982, Trabesinger2012, Georgescu2014]. We identify in particular some families of pyrochlore compounds as suitable frameworks to experimentally test our predictions and suggest how it may be possible to manifest the elusive RVB state for the first time in these materials.
I Results
We study the $t$-$J$ model on a lattice of corner-sharing tetrahedra [footnote:checkerboard] (see Fig. 1) near half-filling, with Hamiltonian
^$\hat{\mathcal{H}}=-t\sum_{\langle i,j\rangle\;\sigma}\left[\hat{c}^{\dagger}_{i\sigma}\hat{c}_{j\sigma}+h.c.\right]+J\sum_{\langle i,j\rangle}\hat{\bm{S}}_{i}\cdot\hat{\bm{S}}_{j}\,,$ (1)
where $\langle i,j\rangle$ denotes nearest-neighbour pairs of sites; $\hat{c}^{\dagger}_{i}$ ($\hat{c}_{i}$) is the constrained fermionic creation (annihilation) operator at site $i$, acting in the space where double occupancy is strictly forbidden (see Supp. Mat. Note C); and $\hat{\bm{S}}_{i}$ is the corresponding spin operator, $\hat{\bm{S}}_{i}=\sum_{\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{i\sigma}{\bm{\sigma}}_{\sigma\sigma^{\prime}}\hat{c}_{i\sigma^{\prime}}$. Here $\sigma$ labels the states of spin-$1/2$ degrees of freedom, and ${\bm{\sigma}}$ is the conventional vector of Pauli matrices. For generality, we treat the hopping $t$ and the spin exchange $J$ as independent parameters. Note that for fermions and our convention in Eq. (1), hole doping with $t>0$ or electron doping with $t<0$ correspond to frustrated hopping when the elementary loop of the lattice has an odd number of sites. In the following we shall focus, without loss of generality, on the case of hole doping and $t>0$.
Our first result is analytical, derived in the thermodynamic limit for a single hole and $J=0$. As demonstrated in detail in the Supp. Mat. Note A, the spectrum of the Hamiltonian in Eq. (1) has a strict lower bound of $-4t$. Furthermore, we prove that this bound is met by the RVB spin liquid ground state
$|\Psi_{\rm RVB}\rangle\propto\sum_{{\bm{r}},\{a_{\bm{r}}\}}\prod_{\langle i,j\rangle\in a_{\bm{r}}}\hat{d}^{\dagger}_{ij}|0\rangle\,,$ (2)
where ${\bm{r}}$ spans all possible positions of the holon; $a_{\bm{r}}$ is a generic dimer covering of the lattice with a holon at ${\bm{r}}$ subject to the condition that there is one and only one dimer per tetrahedron (see e.g., Fig. 1); and $\hat{d}^{\dagger}_{ij}=\frac{1}{\sqrt{2}}(\hat{c}^{\dagger}_{i\uparrow}\hat{c}^{\dagger}_{j\downarrow}-\hat{c}^{\dagger}_{i\downarrow}\hat{c}^{\dagger}_{j\uparrow})$. We shall refer to each state in the superposition forming $|\Psi_{\rm RVB}\rangle$ as a dimer-singlet state. The kinetic energy frustration disappears for the opposite sign of the hopping amplitude, which is equivalent to inserting a $\pi$-flux on each triangular plaquette of the pyrochlore lattice. The RVB wave function minimises the holon kinetic energy by spontaneously generating a $\pi$-flux on the triangular face containing the hole and a singlet. Namely, the RVB wave function binds two $\pi$-fluxes (fluctuating with the dimer-singlet resonant states) to each hole to lower its kinetic energy to $-4t$ by partially lifting the frustration (see Fig. 1).
This is an instance of the so-called counter-Nagaoka effect [Haerter2005], whereby frustrated quantum hole hopping leads to antiferromagnetic correlations in a Mott insulator near half-filling. While the phenomenon has been shown to generically lead to conventionally ordered phases in two dimensions [Sposetti2014, Lisandrini2017, Zhang2018, Morera2021, Kim2023], it had not been thus far investigated in three dimensions, and it is all the more remarkable to see it favour a quantum spin liquid state in lattices of corner-sharing tetrahedra (for a brief review on the counter-Nagaoka effect, see Supp. Mat. Note B).
Figure 2: Top panel: Spin structure factor for the 32-site system with one hole (top half) and two holes (bottom half), for $t=1$ and $J=0$. Accessible momenta are marked by grey pluses. Bottom panel: Blue dots show the holon density for a 16-site and a 32-site system doped with a single hole as a function of distance from the pinned spin [footnote1], indicating that the spinon and holon separate. Crosses show the holon-holon correlations when there are two holes present in the system, and they are numerically identical to the holon-spinon correlations.
The ground state Eq. (2) is exact in the thermodynamic limit, where such dimer coverings with a single monomer are well-defined. In a finite lattice of corner-sharing tetrahedra with periodic boundary conditions, introducing a hole necessarily requires an unpaired spin. As expected for an RVB ground state, the spin and the charge of the fermion added to the Mott insulator separate, i.e., the probability of finding the holon within a finite distance from the spinon is zero in the thermodynamic limit. The spinon must necessarily occupy a passive site of the tetrahedron, implying that the holon will never visit the site occupied by the unpaired spin. In other words, the spinon is expected to have infinite mass (static spinon) and the spinon-holon interaction is expected to be repulsive. Both observations are supported by our numerics, which give our second set of results and attest the stability of our RVB liquid phase in finite-size systems, finite hole density, and the presence of interactions.
We studied the behaviour of Eq. (1) in finite systems using exact diagonalisation (ED) on a 16-site cubic cell, and density matrix renormalisation group (DMRG) [White1992, Schollwock2005, Schollwock2011] on a 32-site system, with periodic boundary conditions (for details of the study, see Supp. Mat. Note D). Remarkably, we find that the ground state energy bound $-4t$ is met (when $J=0$) for all system sizes, within numerical precision. A non-trivial degeneracy is observed ($18$ states for 16 sites and $34$ states for 32 sites). Consistently with the analytical considerations above, we find a unique ground state (again with energy $-4t$) when the unpaired spin is pinned to a given lattice site.
Within numerical accuracy, we show that the ground state wave function with a fixed spin is given by the equal amplitude superposition of all dimer-singlet configurations (constrained to no more than one dimer per tetrahedron) with a delocalised holon, for our 16-site and 32-site systems (see Supp. Mat. Note D). This allows us to propose an exact wave function for the single-hole case for all system sizes – which is a highly unusual feat. Such a state has vanishing expectation value of the projector onto the maximum angular momentum for each tetrahedron with no hole [Kivelson_unpub, Batista2004, Raman2005], which we verified to be the case (Supp. Mat. Note D).
As illustrated in Fig. 2, the holon density shows weak repulsive dependence on the distance from the spinon in finite size systems, in agreement with spin-charge separation in the thermodynamic limit, argued earlier on analytical grounds (see also Supp. Mat. Note D). In the top panel of Fig. 2, we show the spin structure factor of our numerical ground states. Characteristic differences that could be measured experimentally seemingly distinguish it from, for example, the classical or quantum Heisenberg spin liquid structure factors (see Supp. Mat. Note D); however future work on larger system sizes is needed to make more precise statements about this. On general grounds, we expect exponentially decaying spin correlations; the nature of the singlet correlations is less clear, as the correspondence with spin ice states [Baskaran1988, Read1989, Fradkin1990, Raman2005] could induce power law behaviour and an emergent $U(1)$ gauge symmetry with gapless photon excitations.
[…]
Intriguingly, the opposite (stoichiometric) limit of vanishing hole density is singular for the 16-site and 32-site systems in question. The wave function given by the equal amplitude superposition of all dimer-singlet states with one and only one dimer per tetrahedron vanishes identically [footnote:vanishing] (see Supp. Mat. Note A). The RVB liquid phase appears to exist solely in presence of hole doping (at least within the remit of our accessible system sizes and periodic boundary conditions), and therefore it can only be stabilised by kinetic energy frustration.
Finally, we turn to look at the concomitant effects of $t$ and $J$ in the system. We consider the case of a 16-site system with a single hole; we fix $t=1$ and vary $J$, spanning both positive and negative values. In Fig. 3, we plot the behaviour of the ground state spin structure factor for all distinct points in reciprocal space. We find that the spin liquid behaviour persists down to a finite negative value of $(J/t)_{c}\simeq-0.052$ (for holon density $6.25$%), suggesting that a transition from ferromagnet to quantum spin liquid (QSL) may be induced on the pyrochlore lattice by frustrated hole doping. The effective doping-induced antiferromagnetic interaction between the spins has been argued to be proportional to the holon density at sufficiently small concentrations [Haerter2005], and we expect our results to correspondingly become more pronounced [footnote:caveat]. Indeed, adding a second hole to the 16-site system, we find $(J/t)_{c}\simeq-0.173$ (holon density $12.5$%), see Supp. Mat. Note E. In Fig. 3, we further compare for reference the case of electron doping, which is not frustrated for $t>0$ and exhibits a kinetic tendency towards ferromagnetic order through the Nagaoka effect.
Figure 3: Left panel: Dependence of the ground state spin structure factor values (for all individual reciprocal lattice points) on $J/t$ ($t>0$) for a 16-site system with one hole. The region where the correlations are antiferromagnetic in spite of ferromagnetic $J$ is shaded. Right panel: same data when the system is doped with one electron, for comparison.
II Discussion
Our results provide a simple and elegant solution to a problem that has eluded researchers for nearly half a century. Through exact results and detailed numerical simulations, we demonstrate that the long-sought-after RVB spin liquid state exhibiting spin-charge separation emerges naturally as ground state of a large-$U$ Hubbard Hamiltonian on lattices of corner-sharing tetrahedra. The key aspect of this emergent behaviour is the kinetic energy frustration of carriers that result from doping the Mott state.
The non-Fermi-liquid behaviour (e.g., spin-charge separation) of such an RVB liquid state implies that the large-$U$ limit of the pyrochlore Hubbard model cannot be smoothly connected to its weak-coupling counterpart (small $U$). Hence, an effective description at large $U$, which is non-perturbative in nature, is essential to fully characterise it. It is interesting to contrast it to the Fermi liquid behaviour argued to occur in spin liquid states stabilised by frustration in lattices of corner-sharing tetrahedra near half-filling in finite-$U$ Hubbard models (see e.g., Refs. Fujimoto2001, Pan2023). This raises the intriguing possibility of a crossover between the two regimes as a function of doping.
The exact results found in this work about the ground state wave function of the system, both in the thermodynamic limit as well as for finite size, raise interesting questions about the nature of the excitations, and in particular the statistics of the holons, which shall be left to future work.
The eminently neat and realistic nature of the Hamiltonian opens the door to further theoretical investigations into the physics of RVB phases, possibly also with the help of implementations on synthetic quantum platforms [Trabesinger2012, Georgescu2014] (note in particular how the leading RVB liquid behaviour is already apparent in relatively small systems of 16 sites). The rare availability of a seemingly exact single-hole ground state wave function for systems of any size provides uncommon ability to compare theory and simulator experiments.
[…]
Our results apply to corner-sharing tetrahedral lattices in general, including the two-dimensional checkerboard lattice. While less common than the pyrochlore lattice in real materials, one wonders whether such geometry could emerge in quasi-2D square lattice compounds, possibly mediated by inter-layer atoms; or perhaps in 2D materials and synthetic quantum platforms. From a theoretical perspective, it would be interesting to investigate in greater detail the robustness of the RVB liquid behaviour and spin-charge separation to an asymmetry between the nearest-neighbour and next-nearest-neighbour (diagonal) hopping terms on the checkerboard lattice.
As speculated earlier, the RVB singlet state in our system could exhibit an emergent $U(1)$ gauge symmetry with power-law singlet correlators and gapless (photon) excitations (in 3D) [Baskaran1988, Read1989, Fradkin1990]. In presence of itinerant (fermionic [Read1989b]) holons, this is reminiscent of models used to investigate Fermi Liquid Star (FL*) behaviour [Senthil2003], and may provide an additional experimental avenue to explore it.
In closing, our work proposes kinetic energy frustration of doped Mott insulators as a novel route to realising robust topologically ordered states, in contrast to the conventional approach based on geometrically frustrated interactions. This holds great promise for the theoretical and experimental discovery of new emergent states of matter. As for a potential connection between kinetically frustrated RVB liquids and superconductivity [Anderson1987], this shall remain an open question for future work.
III Online Methods
We studied a 16-site system consisting of a single conventional cubic unit cell of the pyrochlore lattice (see Supp. Mat. Note C for illustrations) using ED with Lanczos to compute the lowest eigenvalues and corresponding eigenstates in each magnetisation sector. We also considered a 32-site system consisting of a $2\times 2\times 2$ cell constructed from the fcc primitive lattice vectors. For this system size, we used two-site DMRG with an energy convergence criterion of $10^{-10}$ and a maximum bond dimension of $12800$. We imposed periodic boundary conditions, creating a 1D snake path of the system following the numbering shown in the Supp. Mat. Note C. The choice of snake path seems to have little impact on the convergence for our highly entangled 3D system. The DMRG calculations were performed using the TeNPy Library (version 0.10.0) [tenpy].
A single DMRG calculation gives one ground state. When the ground state is degenerate, we biased each ground state and compute correlations as an average over all ground states. In the case of pure hopping ($J=0$), the full ground state manifold was found by pinning the unpaired spin by adding an on-site magnetic field. This gives an overcomplete set of ground states, and a linearly independent set of ground states is then found via Gram-Schmidt decomposition.
A similar approach was deployed to study our system on the checkerboard lattice, using ED for 16 sites (1 and 2 holes) and 24 sites (1 hole), and DMRG for 24 sites (2 holes).
Acknowledgements
We would like to thank J.Chalker, V.Oganesyan, S.Powell, and O.F.Syljuåsen for useful discussions. C.G. acknowledges support from the Aker Scholarship. The work of A.Š. is supported by the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie Actions Grant agreement No. 101104378. D.P. would like to acknowledge the Oxford-ShanghaiTech Collaboration Project for financial support. C.D.B. acknowledges support from the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Award Number DE-SC0022311. This work was supported in part by the Engineering and Physical Sciences Research Council (EPSRC) grants No. EP/T028580/1 and No. EP/V062654/1 (C.C.). For the purpose of open access, the authors have applied a Creative Commons Attribution (CC-BY) licence to any Author Accepted Manuscript version arising from this submission.
Author contributions
The project was conceived by C.C. and C.G., and all authors contributed to its development, with specific focus by C.G. on the numerical efforts; by C.D.B. on the analytical proofs; and by P.A.G. and D.P. on the experimental component. All authors contributed to writing the manuscript.
Competing interests
The authors declare to have no competing interests in this work.
Materials and Correspondence
Correspondence should be addressed to C.C., and queries about the numerical results should be addressed to C.G.
Supplementary Material for:
“Spin-charge separation and resonant valence bond spin liquid in a frustrated doped Mott insulator”
Cecilie Glittum,1 Antonio Štrkalj,1,2 Dharmalingam Prabhakaran,3 Paul A. Goddard,4 Cristian D. Batista,5,6 and Claudio Castelnovo1
1T.C.M. Group, Cavendish Laboratory, JJ Thomson Avenue, Cambridge CB3 0HE, United Kingdom
2Department of Physics, Faculty of Science, University of Zagreb, Bijenička c. 32, 10000 Zagreb, Croatia
3Department of Physics, University of Oxford, Clarendon Laboratory, Parks Road, Oxford OX1 3PU, UK
4Department of Physics, University of Warwick, Coventry CV4 7AL, United Kingdom
5Department of Physics and Astronomy, University of Tennessee, Knoxville, TN 37996, USA
6Neutron Scattering Division, Oak Ridge National Laboratory, Oak Ridge, TN 37831, USA
SM Note A Proofs
In this section, we present the proof of the first result in our work, relating to the ground state (GS) of Eq. (1) in the main text, for a single hole in the thermodynamic limit when $J=0$ and $t>0$ (frustrated hopping). We note that the proof applies generally to lattices of corner-sharing tetrahedra, including the three-dimensional pyrochlore lattice, as well as the two-dimensional checkerboard lattice.
A.1 Ground state energy
Since the Hamiltonian $\hat{\mathcal{H}}$ includes only nearest neighbour hopping terms, it can be decomposed as a sum of Hamiltonians ^$\hat{\mathcal{H}}^{\alpha}$ acting on each tetrahedron $\alpha$:
^$\hat{\mathcal{H}}=\sum_{\alpha}\hat{\mathcal{H}}^{\alpha}\,,$ (3)
where the dual lattice index $\alpha$ labels the different tetrahedra and ^$\hat{\mathcal{H}}^{\alpha}$ includes only the six bonds of the tetrahedron $\alpha$.
The ground state $|\Psi_{0}\rangle$ can be expressed as
$|\Psi_{0}\rangle=\sum_{{\bm{r}}}|\phi_{\bm{r}}\rangle\,,$ (4)
where $|\phi_{\bm{r}}\rangle$ represents a state where the holon occupies the site ${\bm{r}}$ and the rest of the spins are in an arbitrary spin configuration, $\langle\phi_{\bm{r}}|\phi_{{\bm{r}}^{\prime}}\rangle=\delta_{{\bm{r}},{\bm{r}}^{\prime}}\langle\phi_{\bm{r}}|\phi_{{\bm{r}}}\rangle$, and $\langle\Psi_{0}|\Psi_{0}\rangle=1$.
After introducing the states
$|\chi_{\alpha}\rangle=\sum_{{\bm{r}}\in\alpha}|\phi_{\bm{r}}\rangle\,,$ (5)
where the sum runs over the $4$ sites of the tetrahedron $\alpha$, we compute the expectation value of the energy,
$E_{g}=\sum_{{\bm{r}},{\bm{r}}^{\prime}}\langle\phi_{\bm{r}}|\hat{\mathcal{H}}|\phi_{\bm{r}^{\prime}}\rangle=\sum_{\langle{\bm{r}},{\bm{r}}^{\prime}\rangle}\langle\phi_{\bm{r}}|\hat{\mathcal{H}}|\phi_{\bm{r}^{\prime}}\rangle=\sum_{\alpha}\langle\chi_{\alpha}|\hat{\mathcal{H}}^{\alpha}|\chi_{\alpha}\rangle\,,$ (6)
where $\langle{\bm{r}},{\bm{r}}^{\prime}\rangle$ restricts the sum to nearest-neighbour sites. By introducing the probability distribution over tetrahedra:
$p_{\alpha}=\frac{1}{2}\langle\chi_{\alpha}|\chi_{\alpha}\rangle\,,$ (7)
where the normalisation condition $\sum_{\alpha}p_{\alpha}=1$ follows from $\langle\Psi_{0}|\Psi_{0}\rangle=1$, we express $E_{g}$ as
$E_{g}=\sum_{\alpha}p_{\alpha}2\frac{\langle\chi_{\alpha}|\hat{\mathcal{H}}|\chi_{\alpha}\rangle}{\langle\chi_{\alpha}|\chi_{\alpha}\rangle}\,.$ (8)
Now, assume that $E_{g}<-4t$. Since $p_{\alpha}$ is a probability distribution over tetrahedra, this assumption implies that there is at least one tetrahedron $\beta$ for which
$2\frac{\langle\chi_{\beta}|\hat{\mathcal{H}}|\chi_{\beta}\rangle}{\langle\chi_{\beta}|\chi_{\beta}\rangle}<-4t\quad{\rm or}\quad\frac{\langle\chi_{\beta}|\hat{\mathcal{H}}|\chi_{\beta}\rangle}{\langle\chi_{\beta}|\chi_{\beta}\rangle}<-2t\,.$ (9)
Eq. (9) implies that the ground state energy of a single tetrahedron is lower than $-2t$. However, the ground state energy of a single tetrahedron is exactly $-2t$, which implies that $E_{g}$ cannot be lower than $-4t$.
For completeness, let us close by showing explicitly that the single tetrahedron GS energy is $-2t$. If we consider a system consisting of four fully connected sites, labelled $1,2,3,4$, with one holon and three spins, we can define the wave functions
$|\Psi^{\rm tet}_{1\sigma}\rangle$ $=$ $\frac{1}{\sqrt{3}}\hat{c}^{\dagger}_{1\sigma}\left(\hat{d}^{\dagger}_{23}+\hat{d}^{\dagger}_{34}+\hat{d}^{\dagger}_{42}\right)|0\rangle$
$|\Psi^{\rm tet}_{2\sigma}\rangle$ $=$ $\frac{1}{\sqrt{3}}\hat{c}^{\dagger}_{2\sigma}\left(\hat{d}^{\dagger}_{13}+\hat{d}^{\dagger}_{34}+\hat{d}^{\dagger}_{41}\right)|0\rangle$
$|\Psi^{\rm tet}_{3\sigma}\rangle$ $=$ $\frac{1}{\sqrt{3}}\hat{c}^{\dagger}_{3\sigma}\left(\hat{d}^{\dagger}_{12}+\hat{d}^{\dagger}_{24}+\hat{d}^{\dagger}_{41}\right)|0\rangle\,,$ (10)
with $\hat{d}^{\dagger}_{ij}=\frac{1}{\sqrt{2}}(\hat{c}^{\dagger}_{i\uparrow}\hat{c}^{\dagger}_{j\downarrow}-\hat{c}^{\dagger}_{i\downarrow}\hat{c}^{\dagger}_{j\uparrow})$. Note that $\hat{d}^{\dagger}_{ij}=\hat{d}^{\dagger}_{ji}$. The state $|\Psi^{\rm tet}_{i\sigma}\rangle$ has a static spin $\sigma$ on site $i$, while the holon and a singlet resonate on the remaining three sites. Consequently, the total spin of the state is $\sigma$. Applying the single tetrahedron Hamiltonian to this state, we find that hopping to site $i$ indeed is cancelled, and the states are eigenstates with energy $-2t$. Note that the state where site $4$ has a static spin is a superposition of the three states in Eq. (10): $|\Psi^{\rm tet}_{4\sigma}\rangle=\frac{1}{\sqrt{3}}\hat{c}^{\dagger}_{4\sigma}\left(\hat{d}^{\dagger}_{12}+\hat{d}^{\dagger}_{23}+\hat{d}^{\dagger}_{31}\right)|0\rangle=-(|\Psi^{\rm tet}_{1\sigma}\rangle+|\Psi^{\rm tet}_{2\sigma}\rangle+|\Psi^{\rm tet}_{3\sigma}\rangle)$.
A.2 Ground state wave function
Consider now the resonant valence bond (RVB) state:
$|\Psi_{\rm RVB}\rangle=\frac{1}{\sqrt{\kappa N}}\sum_{{\bm{r}}}|\phi^{\rm RVB}_{\bm{r}}\rangle\,,$ (11)
with
$|\phi^{\rm RVB}_{\bm{r}}\rangle=\sum_{a_{\bm{r}}}|\Psi_{a_{\bm{r}}}\rangle,\quad\kappa=\langle\phi^{\rm RVB}_{\bm{r}}|\phi^{\rm RVB}_{\bm{r}}\rangle\,,$ (12)
where $N\to\infty$ is the number of sites, the index $a_{\bm{r}}$ runs over all dimer coverings with one dimer per tetrahedron on the lattice with the site ${\bm{r}}$ excluded and $|\Psi_{a_{\bm{r}}}\rangle$ is the corresponding direct product of singlet states on each dimer:
$|\Psi_{a_{\bm{r}}}\rangle=\prod_{\langle i,j\rangle\in a_{\bm{r}}}\hat{d}^{\dagger}_{ij}|0\rangle\,.$ (13)
Substituting Eq. (13) into Eq. (12), and thence into Eq. (11), we obtain
$|\Psi_{\rm RVB}\rangle=\frac{1}{\sqrt{\kappa N}}\sum_{{\bm{r}},\{a_{\bm{r}}\}}|\Psi_{a_{\bm{r}}}\rangle=\frac{1}{\sqrt{\kappa N}}\sum_{{\bm{r}},\{a_{\bm{r}}\}}\prod_{\langle i,j\rangle\in a_{\bm{r}}}\hat{d}^{\dagger}_{ij}|0\rangle\,.$ (14)
From Eq. (6), we see that
$E_{\rm RVB}\equiv\langle\Psi_{\rm RVB}|\hat{\mathcal{H}}|\Psi_{\rm RVB}\rangle=\frac{1}{\kappa N}\sum_{\alpha}\langle\chi^{\rm RVB}_{\alpha}|\hat{\mathcal{H}}^{\alpha}|\chi^{\rm RVB}_{\alpha}\rangle,$ (15)
with
$|\chi^{\rm RVB}_{\alpha}\rangle=\sum_{{\bm{r}}\in\alpha}|\phi^{\rm RVB}_{\bm{r}}\rangle\,.$ (16)
According to Eq. (12),
$|\phi^{\rm RVB}_{\bm{r}}\rangle=|\tilde{\Psi}^{{\bm{r}}_{1}{\bm{r}}_{2}}_{{\bm{r}}}\rangle+|\tilde{\Psi}^{{\bm{r}}_{2}{\bm{r}}_{3}}_{{\bm{r}}}\rangle+|\tilde{\Psi}^{{\bm{r}}_{1}{\bm{r}}_{3}}_{{\bm{r}}}\rangle\,,$ (17)
where ${\bm{r}}_{1},{\bm{r}}_{2},{\bm{r}}_{3}$ denote the $3$ sites of the tetrahedron $\alpha$ that are not occupied by the holon (the pair ${\bm{r}}_{i}{\bm{r}}_{j}$ is equivalent to ${\bm{r}}_{j}{\bm{r}}_{i}$) and
$|\tilde{\Psi}^{{\bm{r}}_{i}{\bm{r}}_{j}}_{{\bm{r}}}\rangle=\sum_{a_{\bm{r}}/d\in({\bm{r}}_{i},{\bm{r}}_{j})}|\Psi_{a_{\bm{r}}}\rangle\,,$ (18)
where the sum runs over dimer coverings $|\Psi_{a_{\bm{r}}}\rangle$ with one holon at site ${\bm{r}}$ and one singlet on the bond $({\bm{r}}_{i}{\bm{r}}_{j})$ with $1\leq i,j\leq 3$ and $i<j$. By inserting Eq. (17) in Eq. (16), we obtain:
$|\chi^{\rm RVB}_{\alpha}\rangle=\sum_{{\bm{r}}\in\alpha,i<j}|\tilde{\Psi}^{{\bm{r}}_{i}{\bm{r}}_{j}}_{{\bm{r}}}\rangle\,.$ (19)
The sum contains $12$ terms corresponding to the $12$ different ways of allocating one holon (in one of the $4$ sites of the tetrahedron) and one singlet in one of the $3$ pairs that can be formed with the $3$ remaining sites. We will group these $12$ terms in $4$ groups, where each group consists of the sum of three states:
$|\tilde{\Psi}^{\alpha}_{{\bm{r}}}\rangle=|\tilde{\Psi}^{{\bm{r}}_{i}{\bm{r}}_{j}}_{{\bm{r}}_{k}}\rangle+|\tilde{\Psi}^{{\bm{r}}_{k}{\bm{r}}_{i}}_{{\bm{r}}_{j}}\rangle+|\tilde{\Psi}^{{\bm{r}}_{j}{\bm{r}}_{k}}_{{\bm{r}}_{i}}\rangle\,.$ (20)
The coordinates ${\bm{r}}_{i},{\bm{r}}_{j},{\bm{r}}_{k}$ denote the $3$ distinct sites of the tetrahedron (one containing the holon and the other two forming a singlet state) that are different from the ‘passive’ site ${\bm{r}}$. We note that when considering only the four sites on the tetrahedron $\alpha$, $|\tilde{\Psi}^{\alpha}_{{\bm{r}}}\rangle$ correspond exactly to the eigenstates of the single tetrahedron where one site is passive and the three remaining sites contain a holon and singlet resonating.
Next, if we denote the coordinates of the $4$ sites of the tetrahedron $\alpha$ by ${\bm{r}}_{a}$ with $1\leq a\leq 4$, we have:
$|\chi^{\rm RVB}_{\alpha}\rangle=|\tilde{\Psi}^{\alpha}_{{\bm{r}}_{1}}\rangle+|\tilde{\Psi}^{\alpha}_{{\bm{r}}_{2}}\rangle+|\tilde{\Psi}^{\alpha}_{{\bm{r}}_{3}}\rangle+|\tilde{\Psi}^{\alpha}_{{\bm{r}}_{4}}\rangle\,,$ (21)
and
$E_{\rm RVB}$ $=$ $\frac{1}{\kappa N}\sum_{\alpha}\langle\chi^{\rm RVB}_{\alpha}|\hat{\mathcal{H}}^{\alpha}|\chi^{\rm RVB}_{\alpha}\rangle$ (22)
$=$ $\frac{1}{\kappa N}\sum_{\alpha}\sum_{a,b=1}^{4}\langle\tilde{\Psi}^{\alpha}_{{\bm{r}}_{a}}|\hat{\mathcal{H}}^{\alpha}|\tilde{\Psi}^{\alpha}_{{\bm{r}}_{b}}\rangle\,.$
Finally, we note that
^$\hat{\mathcal{H}}^{\alpha}|\tilde{\Psi}^{\alpha}_{{\bm{r}}_{a}}\rangle=-2t\,|\tilde{\Psi}^{\alpha}_{{\bm{r}}_{a}}\rangle\,,$ (23)
implying that
$E_{\rm RVB}$ $=$ $-\frac{2t}{\kappa N}\sum_{\alpha}\sum_{a,b=1}^{4}\langle\tilde{\Psi}^{\alpha}_{{\bm{r}}_{a}}|\tilde{\Psi}^{\alpha}_{{\bm{r}}_{b}}\rangle$ (24)
$=$ $-\frac{2t}{\kappa N}\sum_{\alpha}\langle\chi^{\rm RVB}_{\alpha}|\chi^{\rm RVB}_{\alpha}\rangle$
$=$ $-\frac{4t}{\kappa N}\sum_{\bm{r}}\langle\phi^{\rm RVB}_{\bm{r}}|\phi^{\rm RVB}_{\bm{r}}\rangle=-4t\,.$
Since we have shown that $-4t$ is a strict lower bound for the energy, Eq. (24) implies that $|\Psi_{\rm RVB}\rangle$ is an exact ground state of $\hat{\mathcal{H}}$. Eq. (23) clarifies the name ‘passive’ for the site ${\bm{r}}$: since $|\tilde{\Psi}^{\alpha}_{\bm{r}}\rangle$ is an eigenstate of ^$\hat{\mathcal{H}}^{\alpha}$, the holon never visits that site because of the destructive interference produced by the linear superposition (20).
A.3 Considerations about the thermodynamic limit
The proof holds in the thermodynamic limit, as only in that limit can we accommodate one dimer in each tetrahedron. Inserting a hole in a finite lattice with an integer number of tetrahedra leads to an unpaired spin that can be associated with a spinon. In this scenario, the dimer coverings are not complete because there is one tetrahedron which does not contain a dimer. We can always choose that tetrahedron to be located at the boundary of the finite lattice and also to be the one containing the unpaired spin. Since the boundary disappears upon taking the thermodynamic limit, the unpaired spin and the missing dimer become irrelevant in this limit. However, this simple observation has two important implications. Firstly, as expected for an RVB ground state, the spin and the charge of the fermion added to the Mott insulator separate, i.e., the probability of finding the holon at any finite distance from the spinon is zero in the thermodynamic limit. Secondly, on a finite size lattice the spinon must necessarily occupy a passive site of the tetrahedron, implying that the holon will never visit the site occupied by the unpaired spin. In other words, the spinon is expected to have infinite mass (static spinon) and the spinon-holon interaction is expected to be repulsive. Both observations are supported by our numerical results.
A.4 Proposed finite-size ground state wave function
From numerics (discussed in Supp. Mat. Notes C and D), we find that the one-hole GS energy $-4t$ can be met in finite size systems and even when the unpaired spin is fixed at a given site; in the latter case the GS is unique. This allows us to conjecture a form for the wave function that is given by the equal amplitude superposition of all dimer-singlet states that have a spin fixed and a delocalised holon, subject to the constraint of no more than one dimer per tetrahedron. Remarkably, we find that the conjectured state is identical within numerical accuracy to the ground state found for the 16-site system using ED, and to the 32-site system using DMRG. We also find this to be the case for the two-holes ground state (i.e., the equal amplitude superposition of all dimer-singlet states that have two delocalised holons, subject to the constraint of no more than one dimer per tetrahedron). We therefore propose that such GS wave functions are in fact exact for all system sizes.
Note that any dimer configuration with a fixed spin and a hole (or two holes) necessarily has one less dimer than tetrahedra. Under the condition of no more than one dimer per tetrahedron, this means that there is precisely one tetrahedron without dimers in the configuration. Notably, we find that the proposed GS wave function is invariant to fixing the location of this tetrahedron.
A.5 Vanishing wave function in the stoichiometric limit
Curiously, we find that our problem does not admit a straightforward stoichiometric limit of vanishing hole density. If we take the equal amplitude superposition of all dimer-singlet states on the pyrochlore lattice, without holons or unpaired spins, the resulting wave function vanishes identically for both the 16-site and the 32-site system. This surprising result – which to the best of our knowledge had not been pointed out in the literature before – holds irrespective of whether we impose the constraint of one and only one dimer per tetrahedron (relevant to our RVB liquid state) or we consider all possible dimer configurations in general [footnote:vanishing2].
SM Note B The counter-Nagaoka effect
Itinerant holons in large-$U$ Hubbard models near half-filling are known to influence the spin correlations. For example, Nagaoka’s theorem [Nagaoka1966, Thouless1965] shows that a single hole in the absence of interactions between the spins is capable of inducing a ferromagnetic ground state. The phenomenon is rooted in the fact that spin correlations affect the kinetic energy of the holon. In Nagaoka’s case, a ferromagnetic pattern allows for perfect constructive interference of the holon world lines, thus minimising the kinetic energy. The effect is remarkably stable (namely, a finite magnetisation persists) up to 20% hole doping [Liang1995, Park2008, Carleo2011, Liu2012] (however, see also Ref. Eisenberg2002 for considerations about a possible breakdown mechanism due to phase separation).
More recently, Haerter and Shastry [Haerter2005] pointed out that the same kinetic mechanism can give rise to strikingly different spin correlations – namely, weak metallic antiferromagnetism – if the electronic lattice is frustrated; this occurs, for instance, when the electron hopping amplitude is negative, and the elementary loops on the lattice encompass an odd number of bonds. This counterpart to the Nagaoka effect was further studied in Refs. Sposetti2014, Lisandrini2017 and later dubbed the ‘counter-Nagaoka effect’ [Zhang2018].
The nature of the state on the frustrated triangular lattice was investigated numerically [Sposetti2014, Lisandrini2017, Morera2021] and shown to exhibit correlations that depart significantly from the behaviour expected for the quantum Heisenberg antiferromagnet, towards a more classical nature (although not entirely: a small deviation from classical correlations is observed [Sposetti2014]). This work also demonstrated that the kinetic energy mechanism was akin to a finite spin interaction term proportional to the hole density (namely, $J_{t}\sim t/N$, where $N$ is the number of lattice sites, in the presence of a single hole).
The effect has been recently investigated in ladders in Ref. Morera2021; and on a Husimi cactus of triangles (akin locally to a kagome lattice) where it was shown to form a regular valence bond pattern [Kim2023].
To date, studies of the counter-Nagaoka effect have focused on 2D (or quasi-1D) systems, where it is found to lead to magnetically ordered low-energy states. In this work, we explore the behaviour on the 3D pyrochlore lattice (see Fig. 1 in the main text), renowned for its strong frustration effects (and our results straightforwardly apply also to a planar arrangement of tetrahedra, also known as the 2D checkerboard lattice).
SM Note C Models and methods
Figure S1: 16-site pyrochlore system, constructed from a single cubic unit cell with lattice vectors ${\bm{a}}_{1}=(1,0,0)$, ${\bm{a}}_{2}=(0,1,0)$ and ${\bm{a}}_{3}=(0,0,1)$.
Figure S2: 32-site pyrochlore system, constructed from a $2\times 2\times 2$ lattice of the ffc unit cell with lattice vectors ${\bm{a}}_{1}=(0,1/2,1/2)$, ${\bm{a}}_{2}=(1/2,0,1/2)$ and ${\bm{a}}_{3}=(1/2,1/2,0)$. Each unit cell has four sites (a single tetrahedron).
In the main text, we expressed the Hamiltonian for our system in terms of constrained electron operators for notational convenience. The constraint can be made explicit by substituting $\hat{c}_{j\sigma}\to\hat{c}_{j\sigma}\hat{P}_{j\sigma}$, where on the right hand side we have the conventional femionic annihilation operator, and $\hat{P}_{i\sigma}=(1-\hat{n}_{i\overline{\sigma}})$ is the projection operator that ensures no double occupancy, with $\hat{n}_{i\sigma}$ being the customary density operator. This leads to the Hamiltonian written as
^$\hat{\mathcal{H}}=-t\sum_{\langle i,j\rangle\;\sigma}\left[\hat{P}_{i\sigma}\hat{c}^{\dagger}_{i\sigma}\hat{c}_{j\sigma}\hat{P}_{j\sigma}+h.c.\right]+J\sum_{\langle i,j\rangle}\hat{{\bm{S}}}_{i}\cdot\hat{{\bm{S}}}_{j}\,,$ (25)
where once again $\langle i,j\rangle$ denotes nearest-neighbour pairs of sites; $\hat{c}^{\dagger}_{i}$ ($\hat{c}_{i}$) is the fermionic creation (annihilation) operator at site $i$; and $\hat{{\bm{S}}}_{i}$ is the corresponding spin operator, $\hat{{\bm{S}}}_{i}=\sum_{\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{i\sigma}{\bm{\sigma}}_{\sigma\sigma^{\prime}}\hat{c}_{i\sigma^{\prime}}$.
We study 16- and 32-site systems of the pyrochlore lattice, both with periodic boundary conditions. The 16-site system is a conventional cubic unit cell, illustrated in Fig. S1. The 32-site system is a $2\times 2\times 2$ spherical cell constructed from the fcc primitive lattice vectors, illustrated in Fig. S2.
The 16-site and 32-site systems are chosen as they respect the symmetries of the pyrochlore lattice. Other shapes of the unit cell show a tendency to seemingly spurious finite-size effects at very low temperatures in a way akin to partial order-by-disorder effects induced by shape anisotropy. The next possible symmetric system sizes would be the $3\times 3\times 3$ spherical fcc system, with 108 sites, or a $2\times 2\times 2$ cubic system with $128$ sites. However, these system sizes are beyond the computational reach of our study.
For the 16-site system, we perform ED with Lanczos to compute the $800$ smallest eigenvalues and corresponding eigenstates in each magnetisation sector. We ensure we capture all ground states by checking that their number stays the same when the number of computed eigenvalues and eigenstates is doubled.
We use two-site DMRG to study the 32-site system, with an energy convergence criterion of $10^{-10}$ and a maximum bond dimension up to $12800$. Using periodic boundary conditions, we create a 1D snake path of the system that follows the numbering in Fig. S2. The choice of snake path seems to have little impact on the convergence for our highly entangled 3D system. The DMRG calculations were performed using the TeNPy Library (version 0.10.0) [tenpy].
A single DMRG calculation gives one ground state. When the ground state is degenerate due to an odd number of spins, we bias each ground state and compute correlations as an average over all ground states. In the case of pure hopping ($J=0$, $t=1$), the full GS manifold is found by pinning the unpaired spin by adding an on-site magnetic field. This gives an overcomplete set of ground states, and a linearly independent set of ground states is then found via Gram-Schmidt decomposition.
For the studies of the classical Heisenberg model in Supp. Mat. Note D, we use Monte Carlo (MC) simulations with the Metropolis algorithm [Newman2011]. The classical structure factors are computed after equilibrating the system down to $\beta=10^{3}$. In order to compare quantitatively classical spins with quantum spin $S=1/2$ degrees of freedom, we tune the classical spin length to satisfy $S^{2}=3/4$.
As mentioned in the main text, our results hold for lattices of corner sharing tetrahedra and therefore also for the 2D checkerboard lattice. Indeed, the 16-site unit cell of the 2D checkerboard lattice is identical to the one for the pyrochlore lattice. In our work, we further consider a 24-site checkerboard lattice using ED (1 hole) and DMRG (2 holes).
SM Note D Doping and correlations
In the first instance, we consider the case of non-interacting spins ($t=1$, $J=0$) at zero temperature in the presence of a single hole. For all three system sizes, we find that the ground state energy bound is met at $-4t$. For 16 and 32 sites, the GS degeneracy is $18$ and $34$, respectively.
An overall 2-fold degeneracy can be readily understood from the odd number of spin-$1/2$ degrees of freedom in the system. The residual degeneracy is clearly related to the presence of an unpaired spin. Indeed, we find a unique ground state when the unpaired spin is pinned to a given lattice site, resulting in a set of $16$ and $32$ linearly-dependent states that span the GS manifold. A set of orthonormal GS wave functions can then be found using Gram-Schmidt decomposition, for example, giving the 9-fold and 17-fold degeneracy. This is in contrast with the behaviour observed for two holes (see Supp. Mat. Note E), where there are no unpaired spins and the GS is unique. While the specific degeneracy of the single-hole ground state for arbitrary system sizes remains to be understood, our results suggest that the scaling is at most linear in system size.
The holon density shows weak repulsive dependence on the distance from the fixed spin, as illustrated in Fig. 2 in the main text. Correspondingly, the expectation value of the hole hopping operator has a suitably inverse dependence, leading to the exact Hamiltonian eigenvalue $-4t$. This is confirmed by computing the expectation value $\langle\sum_{\sigma,j}\hat{c}^{\dagger}_{i,\sigma}\hat{c}_{j,\sigma}\rangle$, where $j$ runs over all nearest-neighbour sites of $i$. Within numerical accuracy, we find indeed that this expectation value is precisely $4\langle 1-\hat{n}_{i}\rangle$ for all sites $i$, where $\langle 1-\hat{n}_{i}\rangle$ is the holon density at site $i$.
We further checked that, within numerical precision, the expectation value of the projector onto the state with maximum angular momentum for each tetrahedron with no holon,
$Q_{\rm tet}=\left\langle\hat{{\bm{S}}}_{\rm tet}^{2}\left(\hat{{\bm{S}}}_{\rm tet}^{2}-2\right)\left(\sum_{i\in{\rm tet}}\hat{n}_{i}-3\right)\right\rangle\,,$ (26)
vanishes identically in the ground state for all tetrahedra. Here, $\hat{{\bm{S}}}_{\rm tet}=\sum_{i\in{\rm tet}}\hat{{\bm{S}}}_{i}$ is the total spin of the given tetrahedron; and $\sum_{i\in{\rm tet}}\hat{n}_{i}$ takes on the value $4$ or $3$ depending on whether a holon is absent or present on the tetrahedron, respectively.
We find within numerical accuracy that the GS wave function (with a fixed spin) from 16-site ED and 32-site DMRG for the pyrochlore lattice and 16- and 24-site ED for the checkerboard lattice is identical to a state constructed as the equal amplitude superposition of dimer-singlet states with a holon and the further constraint of no more than one dimer per tetrahedron.
We compute the spin structure factor of our RVB liquid state, averaged over all ground states of our system for $t=1$ and $J=0$, as $\mathcal{S}({\bm{q}})=\sum_{i,j}\langle S_{i}^{z}S_{j}^{z}\rangle e^{-i{\bm{q}}\cdot({\bm{r}}_{i}-{\bm{r}}_{j})}/N$, where ${\bm{r}}_{i}$ is the position of site $i$. This is shown in Fig. S3 (upper panels) for the 32-site system, obtained using DMRG. There is indeed no evidence of symmetry breaking. While one can recognise features akin to those of the spin-$1/2$ quantum Heisenberg model (obtained using DMRG; see Fig. S3, middle panels), the doping-induced correlations in the structure factor of the RVB liquid state in our work are clearly distinct in a way that promises to be measurable in experiments. We also compare it for reference to the case of a classical Heisenberg AFM obtained using classical MC simulations (Fig. S3, lower panels). Both the quantum and classical pyrochlore Heisenberg AFM are known to be fully frustrated and in a spin liquid GS [footnote:nematic]. In the case of the pure Heisenberg models we do not introduce any holes and we scale the resulting correlations by $(N-N_{h})/N$, where $N=32$ and $N_{h}=1$ are the number of sites and holes in the RVB state, respectively.
Figure S3: Spin structure factors of the 32-site systems. Top: Doping-induced, $1$ hole ($t=1$ and $J=0$). Middle: Quantum Heisenberg model (QHM). Bottom: Classical Heisenberg model (CHM). Both Heisenberg models have no holes and their correlations are correspondingly scaled by $(N-1)/N$. Accessible momenta are marked by grey pluses.
SM Note E Two holes
We compute the GS in the presence of two holes and study the GS spin and holon correlations for $t=1$ and $J=0$. We find that the GS is unique with energy $-8t$ in the $S^{z}=0$ sector for both the 16-site and 32-site systems, which is exactly twice the GS energy of a single hole.
We verified that, within numerical accuracy, the GS wave function obtained with ED for 16 sites and DMRG for 32 sites is identical to the equal amplitude superposition of all dimer-singlet states hosting two holons, subject to the constraint of no more than one dimer per tetrahedron. (We also confirmed that this result holds for the 24-site checkerboard lattice.)
We find that the doping-induced spin correlations when two holes are present (see Fig. 2 in the main text) are similar to the ones obtained for a single hole. Quantitatively, we compare them using the measure:
$D=\frac{1}{N}\sum_{{\bm{q}}}\left[\vphantom{\sum}\mathcal{S}_{\rm 1-hole}({\bm{q}})-\mathcal{S}_{\rm 2-holes}({\bm{q}})\right]^{2}\,,$ (27)
and we find values $\lesssim 10^{-4}$ and $\lesssim 10^{-5}$ for the 16-site and 32-site case, respectively, when $\mathcal{S}_{\rm 1-hole}$ and $\mathcal{S}_{\rm 2-holes}$ are suitably rescaled by $(N-N_{h})/N$ to account for the different total number of spins.
Figure S4: Dependence of the ground state spin structure factor (for all individual reciprocal lattice points) on $J/t$ ($t>0$) for the 16-site system with two holes. Dashed horizontal lines show the $t=0$ limits, for reference. The shaded region denotes correlations that are antiferromagnetic in spite of a ferromagnetic interaction $J$. The dashed vertical line shows $(J/t)_{c}\simeq-0.173$ for holon density $1/16$.
Next, we study the effect of having a finite $J$ when two holes are present in the 16-site system. In Fig. S4 we show the GS spin structure factor as a function of $J/t$ with $t=1$. We find that the spin liquid behaviour persists down to a finite negative value of $(J/t)_{c}\simeq-0.173$ (hole density $12.5$%). This shows that also in the presence of two holes, quantum spin liquid behaviour can be induced over magnetic ordering by frustrated hole doping.
Figure S5: Holon-holon correlations as a function of $J/t$. $\hat{n}_{i}$ is the spin density operator on site $i$. We measure the correlations between a holon at site 0 and a holon at a site either one or two lattice spacings away. The shaded region shows the range of $J/t$ where the spin correlations are antiferromagnetic in spite of a ferromagnetic interaction $J$.
Finally, we study the holon-holon correlations. The dependence on distance between the two holes is shown in Fig. 2 in the main text, for $t=1$ and $J=0$ and system sizes 16 and 32 (and, consistently with the structure of the wave function, it overlaps with the spinon-holon correlation for a single-hole GS). For the 16-site system, we also study the behaviour in presence of interactions, as a function of $J/t$, in Fig. S5. Note that because of the small system size, we only have access to three distances (with periodic boundary conditions, two sites in this system are either one or two lattice spacings apart). Additionally the second-nearest neighbour and third-nearest neighbour correlators give the same values, and we therefore only have two independent values of the correlator. Within the limitation of the small system size accessed in our study, we observe no sign of holon attraction (rather a small tendency to holon repulsion) in the RVB liquid phase. In the large-$J$ ferromagnetic regime, when the GS becomes partially magnetised, holon attraction is introduced by the holons sharing minority spins [Zhang2018].
SM Note F Experiments
Among the known pyrochlore materials, the platinum group (Pt, Pd, Rh, Ru, Ir, and Os) of compounds, particularly the iridates and ruthenates, have special interest due to their mobile charge carriers and strong spin-orbit coupling [Subramanian, WitczakKrempa2012]. As discussed in the main text, iridate pyrochlore oxides (RE2Ir2O7, RE = rare earth) are especially promising because – while being far from a spin-$1/2$ half-filling Hubbard model description – stoichiometric Ir4+ ions have 5d5 shells with effective $J_{\rm eff}=1/2$ ground state doublets, whereas doping-induced Ir5+ have 5d4 shells and $J_{\rm eff}=0$. In these materials, the iridium ions sit on a lattice of corner-sharing tetrahedra and undergo an antiferromagnetic transition to a long-range ordered state that is characterised by an all-in-all-out (AIAO) arrangement in which the iridium moments point directly into/out of neighbouring tetrahedra [Tomiyasu2012, Donnerer2016, Jacobsen2020]. The ordering temperature of the Ir sublattice depends on Ir–O–Ir bond angles and hence is highly sensitive to the size of the RE ion [Wang2020]. As such, to test our predictions effectively, holes should be introduced while perturbing the IrO6 octahedra as little as possible. In addition, any obfuscating effects of magnetism from the RE sublattice should be avoided. This could be achieved via the synthesis of various doping series in which non-magnetic RE ions (e.g., Y, La, Lu) are successively substituted by an ion of similar size but different charge (e.g., Ca or Sr).
F.1 Measurements
Experimentally, the primary indicator of the onset of spin liquid behaviour in the pyrochlore iridates would be the suppression of the AIAO magnetic order on the Ir sublattice as a function of hole doping. In the materials with non-magnetic RE ions, this can be readily observed in the temperature dependence of the magnetic susceptibility. The onset of AIAO order is typically accompanied by a metal-to-insulator transition [Matsuhira2011], and resistivity measurements may also be used to track its suppression (with the caveat that a separation of the magnetic and resistive transitions is also possible, as recently observed in Sm2Ir2O7 under applied pressure [Coak2024]). Other measurements could be undertaken that provide information on the nature of any observed magnetic ordering, including resonant X-ray scattering, muon spin-relaxation and neutron diffraction [Wang2020, Jacobsen2020, Disseler2012]. Crystal structure determination and careful modelling would also need to be performed to disentangle any effects on the magnetic properties arising from structural changes or from the increase in disorder caused by the introduction of random defects, from those anticipated due to the addition of holes.
As discussed earlier, some studies of this kind have already been performed, including measurements of polycrystalline samples of Y2-xCaxIr2O7 for $x=0$–0.2  [Fukazawa2002, Zhu2014]. Using X-ray photoelectron spectroscopy, Ref. Zhu2014 finds that the amount of Ir5+ in the sample does indeed increase with Ca doping and, at the same time, transport measurements indicate that the conductivity improves. Moreover, a feature in the resistivity that might be associated with the metal-insulator transition seen in other pyrochlore iridates moves to lower temperatures. And temperature-dependent magnetic susceptibility measurements show that the bifurcation of the zero-field and field-cooled traces (which likely arises due to weak ferromagnetic behaviour and is typically associated with the onset of AIAO order in the pyrochlore iridates [Ishikawa2012]), shifts up in temperature with doping. This last result, however, is in contradiction to Ref. Fukazawa2002, which finds that the bifurcation moves to lower temperatures over the same doping range. That study also finds an enhancement of the conductivity as $x$ increases, but no features related to any metal-insulator transition could be discerned. One drawback of these studies is the polycrystalline nature of the samples. Broken symmetry at the surface and the unknown effects of intergrain boundaries in pressed pellets are known to affect the experimental results in these materials, particularly transport measurements, and tend to obscure what is happening in the bulk [Coak2024]. So, while a marked effect on the magnetic and transport properties of pyrochlore iridates has already been observed on doping with holes, measurements of high-quality single crystals are necessary to understand these changes and test the predictions discussed here. This requires further investigation into improving synthesis techniques.
F.2 Crystal growth
[…]
While the pyrochlore iridates are a promising testbed for the theoretical predictions discussed here, some deviations from the ideal $J=1/2$ Heisenberg model should be noted. Firstly, the Ir4+ ions in RE2Ir2O7 are known to be Ising-like, with the spins preferring to point along the local [111] directions, i.e., into or out of the tetrahedra. It is this anisotropy, coupled with the AFM interactions, that gives rise to the AIAO ordered state at low temperatures and any reduction in the Ising-like nature of the spins (e.g., by substitution of a RE ion with a smaller radius) leads to a suppression of the transition temperature. Secondly, the Ir ions in these materials also exhibit a degree of mixing of the $J_{\rm eff}=1/2$ state with the $J_{\rm eff}=3/2$ states due to the local trigonal crystal field (although this is found to be minimised by replacement of the RE ion with some other species). The effects of these deviations from the ideal model are yet to be established, but these materials remain a promising avenue for exploration in this context.